The wavelength of light visible to the human
eye is on the order of 5 × 10−7 m.
Find the frequency of the lightwave if the
speed of light in air is 2.99792 × 108 m/s.
Answer in units of s−1.
1
Expert's answer
2020-02-24T11:03:05-0500

The frequency of the lightwave:


"f=\frac{c}{\lambda}"

"f=\frac{2.99792 \cdot 10^8}{5 \cdot 10^{−7}}=5.99584\cdot 10^{14}s^{-1}"